# Accelerating Training of Recursive Reasoning Models with Curriculum Guided Adaptive Recursion

### v2

@@ V-2 Direct Checkpoint Comparison
To isolate the individual contributions of Progressive Depth Curriculum (PDC) and Hierarchical Supervision Weighting (HSW), we conduct a $2\times 2$ factorial ablation with four configurations trained for 30K epochs. Table V reveals that PDC is the dominant component, providing $2.26\times$ speedup while maintaining comparable accuracy ($85.47\%$ vs baseline $85.14\%$), while HSW is complementary, offering $1.61\times$ speedup through variance reduction but with reduced accuracy ($78.63\%$). Their combination yields $1.71\times$ overall speedup with $82.76\%$ accuracy, demonstrating subadditive interaction where benefits partially overlap rather than compound.
TABLE V: Ablation study summary (30K epochs). Curriculum-Only achieves best speedup ($2.26\times$) and accuracy ($85.47\%$).
@@
    Ablation studies trained for 30K epochs. Main comparison (Table I) used 50K epochs.
@@
Table VI presents detailed metrics including loss decomposition, reasoning efficiency, and halting behavior across all configurations.
TABLE VI: Detailed ablation metrics (30K epochs): losses, reasoning steps, and halting quality.
Config. LM Halt Avg Q-Halt It/sec
Loss Loss Steps Acc(%)
Baseline 0.635 0.025 5.85 98.5 2.36
+ PDC only 0.623 0.014 5.52 100.0 12.05
+ HSW only 0.613 0.030 5.99 98.29 1.37
+ PDC + HSW 0.615 0.020 5.50 99.5 2.84
@@
V-3 Time-to-Accuracy Analysis
Table VII analyzes convergence speed by measuring wall-clock time required to reach specific accuracy milestones, revealing that curriculum dramatically accelerates early learning.
TABLE VII: Time to reach accuracy milestones (from checkpoint evaluations). Curriculum-Only reaches 80% accuracy $1.37\times$ faster than baseline.
Config. To 70% To 75% To 80% To 85%
Baseline 2.5h 3.8h 4.8h 6.6h
+ PDC only 1.5h 2.2h 3.5h 4.5h
+ HSW only 3.5h 5.0h 6.0h N/A
+ Both 1.5h 2.5h 3.0h 5.7h
Speedup to 80% accuracy
[…]
Curriculum-Only reaches the critical 80% threshold in 3.5 hours, $1.37\times$ faster than baseline’s 4.8 hours, and uniquely continues improving to 85.47% in 4.5 hours, a milestone baseline achieves only at 6.6 hours ($1.47\times$ speedup). Combined CGAR achieves even faster convergence to 80% in 3.0 hours ($1.60\times$ faster than baseline). This demonstrates that curriculum provides not only faster initial convergence but also superior final representations that generalize better.
@@ V-3 Time-to-Accuracy Analysis
Figure 4: CGAR generalization analysis across checkpoints. Consistent $\sim$1.3% train-test gap indicates excellent generalization without overfitting.
@@
V-4 Curriculum Phase Transitions
Table VIII tracks Curriculum-Only training progression across the full 30K epoch trajectory, including curriculum phase transitions at $\rho=0.3$ (30% progress, 9K epochs) and $\rho=0.6$ (60% progress, 18K epochs).
TABLE VIII: Curriculum-Only training progression showing phase transitions and learning acceleration.
Epoch Progress Phase Exact Token
2K 7 Shallow 24.3 74.8
4K 13 Shallow 39.7 81.9
6K 20 Shallow 51.2 86.1
8K 27 Shallow 57.4 87.9
10K 33 Medium 64.8 89.7
12K 40 Medium 69.5 90.8
14K 47 Medium 74.2 91.7
16K 53 Medium 77.6 92.6
18K 60 Full 79.8 93.3
20K 67 Full 81.4 93.8
24K 80 Full 83.7 94.3
28K 93 Full 84.9 94.7
30K 100 Full 85.47 94.87
Curriculum phase transitions occur smoothly without accuracy degradation. The Shallow phase (epochs 0-9K, depth $(2,1)$) establishes fundamental constraint satisfaction with rapid early learning: exact accuracy climbs from $24.3\%$ to $57.4\%$ in 8K epochs. The Medium phase (epochs 9K-18K, depth $(4,2)$) refines reasoning capacity, achieving $77.6\%$ by epoch 16K. The Full phase (epochs 18K-30K, depth $(6,3)$) enables maximum reasoning depth, reaching final $85.47\%$ accuracy. Token accuracy exhibits slower growth ($74.8\%$ to $94.87\%$) compared to exact accuracy’s dramatic improvement, consistent with staged learning: early phases learn local prediction rules, later phases integrate them into globally consistent solutions.
The learning acceleration pattern reveals that most accuracy gains occur during shallow and medium phases ($24.3\%$ to $77.6\%$ in first 16K epochs), while the full-depth phase provides final refinement ($79.8\%$ to $85.47\%$ in last 12K epochs). This validates the curriculum design: constrained-depth training prevents overfitting to complex patterns before basics are mastered, producing superior representations that generalize effectively even when full computational capacity becomes available.
@@ V-4 Curriculum Phase Transitions
To validate our choice of hierarchical supervision decay parameter $\lambda=0.7$, we conduct a sensitivity analysis across the range $\lambda\in\{0.5,0.6,0.7,0.8,0.9\}$. Table IX reveals a U-shaped performance curve where $\lambda=0.7$ achieves optimal accuracy ($87.3\%$), while extreme values cause significant issues: $\lambda=0.5$ is too aggressive, leading to training instability and collapse to $22\%$ accuracy, while $\lambda=0.9$-$1.0$ underweight early supervision steps, degrading to $76.8\%$ accuracy. The method demonstrates robustness within $\lambda\in[0.6,0.8]$, providing practitioners with reasonable tuning flexibility.
TABLE IX: Hierarchical supervision decay sensitivity analysis across different exponential weighting schedules. Accuracies measured on training set for controlled comparison.
@@
[…]
Conservative weighting ($\lambda\geq 0.8$) shows graceful degradation: $\lambda=0.8$ achieves respectable $83.10\%$ accuracy (4.22-4.22 points vs optimal), while $\lambda=0.9$ reaches $76.80\%$ (10.52-10.52 points). These configurations learn slower but avoid catastrophic failure all reasoning steps receive relatively uniform supervision, preventing over-specialization on early predictions.
@@
This sensitivity analysis underscores that hierarchical supervision weighting requires careful tuning unlike progressive curriculum, which shows robust performance across reasonable schedules, the exponential decay parameter significantly impacts whether recursive models successfully learn to refine predictions across multiple reasoning steps.
V-A Computational Analysis
@@
$\mathbb{E}_{\rho}[\mathcal{D}_{\text{eff}}]=\sum_{i=1}^{3}\tau_{i}\cdot\mathcal{D}_{\text{eff}}^{(i)}=0.3\cdot 6+0.3\cdot 20+0.4\cdot 42=24.6\text{ layers}$ (7)
@@
VI Conclusion
This paper has presented CGAR (Curriculum-Guided Adaptive Recursion), a training methodology that applies curriculum learning to architectural recursion depth rather than data ordering. CGAR consists of two components: Progressive Depth Curriculum dynamically adjusts recursion parameters $(n,T)$ from shallow to deep configurations during training, while Hierarchical Supervision Weighting applies exponentially decaying importance $w_{t}=\lambda^{t-1}$ to supervision steps based on observed gradient magnitude decay in recursive architectures. Experimental evaluation on Sudoku-Extreme with 423,168 test puzzles demonstrates $1.71\times$ training speedup (10.93 to 6.38 hours) with 0.63% accuracy reduction (86.65% to 86.02%). Systematic ablation studies reveal Progressive Depth Curriculum achieves $2.26\times$ speedup with comparable accuracy (85.47% vs 85.14% baseline), while Hierarchical Supervision Weighting provides $1.61\times$ speedup through 40% gradient variance reduction.
The approach treats architectural depth $\mathcal{D}_{\text{eff}}(n,T)$ as a curriculum-scheduled training parameter rather than a fixed constant. This enables computational savings during training while preventing early-stage overfitting. CGAR-trained models demonstrate improved inference efficiency with 100% halting accuracy and 11% fewer reasoning steps compared to baseline. The 42% training cost reduction improves accessibility for resource-constrained research environments. By reducing training time from 10.93 to 6.38 hours on standard hardware, CGAR makes recursive reasoning models more practical for broader adoption in neurosymbolic AI, program synthesis, and interpretable reasoning systems.
VI-A Limitations and Future Work

### v3

@@ 5.0.2. Checkpoint Comparison at Matching Training Steps
To isolate the individual contributions of Progressive Depth Curriculum (PDC) and Hierarchical Supervision Weighting (HSW), we conduct a $2\times 2$ factorial ablation with four configurations trained for 30K epochs. Table 4 reveals that PDC is the dominant component, providing $2.26\times$ speedup while maintaining comparable accuracy ($85.47\%$ vs baseline $85.14\%$), while HSW is complementary, offering $1.61\times$ speedup through variance reduction but with reduced accuracy ($78.63\%$). Their combination yields $1.71\times$ overall speedup with $82.76\%$ accuracy, demonstrating subadditive interaction where benefits partially overlap rather than compound.
Table 4. Ablation study summary (30K epochs). Curriculum-Only achieves best speedup ($2.26\times$) and accuracy ($85.47\%$).
@@
    Ablation studies trained for 30K epochs. Main comparison (Table 1) used 50K epochs.
@@
Figure 3 presents detailed metrics including loss decomposition, reasoning efficiency and halting behavior across all configurations.
Figure 3. Detailed ablation metrics (30K epochs): losses, reasoning steps and halting quality. Green bars highlight best performance for each metric. PDC-only achieves perfect halting accuracy (100%) with lowest halt loss (0.014) and highest iteration speed (12.05 it/sec), while HSW-only achieves lowest LM loss (0.613).
@@
5.0.3. Time-to-Accuracy Analysis
Figure 4 analyzes convergence speed by measuring wall-clock time required to reach specific accuracy milestones, revealing that curriculum dramatically accelerates early learning.
Figure 4. Time to reach accuracy milestones across training. Two-panel visualization shows Early Convergence (70-80%) and Final Push (80-85%). Curriculum-Only (green) reaches 80% accuracy $1.37\times$ faster than baseline, while combined CGAR (blue) achieves $1.60\times$ speedup. HSW-only (red) does not reach 85% accuracy.
Curriculum-Only reaches the critical 80% threshold in 3.5 hours, $1.37\times$ faster than baseline’s 4.8 hours and uniquely continues improving to 85.47% in 4.5 hours, a milestone baseline achieves only at 6.6 hours ($1.47\times$ speedup). Combined CGAR achieves even faster convergence to 80% in 3.0 hours ($1.60\times$ faster than baseline). This demonstrates that curriculum provides not only faster initial convergence but also superior final representations that generalize better.
@@ 5.0.3. Time-to-Accuracy Analysis
Figure 5. CGAR generalization analysis across checkpoints. Consistent $\sim$1.3% train-test gap indicates excellent generalization without overfitting.
5.0.4. Curriculum Phase Transitions
Fig 6 tracks Curriculum-Only training progression across the full 30K epoch trajectory, including curriculum phase transitions at $\rho=0.3$ (30% progress, 9K epochs) and $\rho=0.6$ (60% progress, 18K epochs).
Figure 6. Curriculum-Only training progression showing phase transitions and learning acceleration
@@ 5.0.4. Curriculum Phase Transitions
To validate our choice of hierarchical supervision decay parameter $\lambda=0.7$, we conduct a sensitivity analysis across the range $\lambda\in\{0.5,0.6,0.7,0.8,0.9\}$. Table 5 reveals a U-shaped performance curve where $\lambda=0.7$ achieves optimal accuracy ($87.3\%$), while extreme values cause significant issues: $\lambda=0.5$ is too aggressive, leading to training instability and collapse to $22\%$ accuracy, while $\lambda=0.9$-$1.0$ underweight early supervision steps, degrading to $76.8\%$ accuracy. The method demonstrates robustness within $\lambda\in[0.6,0.8]$, providing practitioners with reasonable tuning flexibility.
Table 5. Hierarchical supervision decay sensitivity analysis across different exponential weighting schedules. Accuracies measured on training set for controlled comparison.
@@
This sensitivity analysis demonstrates that hierarchical supervision weighting requires careful tuning unlike progressive curriculum, which shows strong performance across reasonable schedules, the exponential decay parameter critically affects whether recursive models successfully learn to refine predictions across multiple reasoning steps.
Figure 7. Decay Sensitivity Analysis: $\lambda=0.5$ vs $\lambda=0.7$
5.1. Computational Analysis
@@
(8) $\mathbb{E}_{\rho}[\mathcal{D}_{\text{eff}}]=\sum_{i=1}^{3}\tau_{i}\cdot\mathcal{D}_{\text{eff}}^{(i)}=0.3\cdot 6+0.3\cdot 20+0.4\cdot 42=24.6\text{ layers}$
@@
6. Conclusion
In this research we presented CGAR (Curriculum-Guided Adaptive Recursion), a training methodology that applies curriculum learning to architectural recursion depth rather than data ordering. CGAR consists of two components: Progressive Depth Curriculum dynamically adjusts recursion parameters $(n,T)$ from shallow to deep configurations during training, while Hierarchical Supervision Weighting applies exponentially decaying importance $w_{t}=\lambda^{t-1}$ to supervision steps based on observed gradient magnitude decay in recursive architectures. Experimental evaluation on Sudoku-Extreme with 423,168 test puzzles demonstrates $1.71\times$ training speedup (10.93 to 6.38 hours) with 0.63% accuracy reduction (86.65% to 86.02%). Systematic ablation studies reveal Progressive Depth Curriculum achieves $2.26\times$ speedup with comparable accuracy (85.47% vs 85.14% baseline), while Hierarchical Supervision Weighting provides $1.61\times$ speedup through 40% gradient variance reduction.
The approach treats architectural depth $\mathcal{D}_{\text{eff}}(n,T)$ as a curriculum-scheduled training parameter rather than a fixed constant. This enables computational savings during training while preventing early-stage overfitting. CGAR-trained models demonstrate improved inference efficiency with 100% halting accuracy and 11% fewer reasoning steps compared to baseline. The 42% training cost reduction improves accessibility for resource-constrained research environments. By reducing training time from 10.93 to 6.38 hours on standard hardware, CGAR makes recursive reasoning models more practical for broader adoption in neurosymbolic AI, program synthesis and interpretable reasoning systems.
7. Limitations and Future Work
@@
Appendix A Benchmark Task Walkthroughs
Figure 8. CGAR solving a Sudoku-Extreme puzzle in 2 reasoning steps, demonstrating adaptive computation with early halting after sufficient reasoning.
Figure 9. Panel view of CGAR on difficult Sudoku puzzles showing adaptive recursion depth $(n,T)$ adjustment based on problem complexity.
Figure 10. CGAR solving medium-difficulty Sudoku puzzles with fewer reasoning iterations, matching computational effort to problem difficulty.